OpenStudy (anonymous):

@iambatman Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of ten inches and a height of 3 inches. Each leg is 3 inches by 3 inches by 12 inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

Ok so find the surface area

OpenStudy (anonymous):

You need to find the surface area of the rectangular prism, \[3 \times 3 \times 12\], mhm it seems we have to make some assumptions as well

OpenStudy (anonymous):

3 x 3 x 12 = 108

OpenStudy (anonymous):

Ok I guess find the surface area of the seat, \[SA = (2 \pi r h) +(2 \pi r^2)\] (surface area of the cylinder) also we need the surface area of the legs \[S = 2lw+2lh+2wh\]

OpenStudy (anonymous):

So now you just have to plug in the numbers

OpenStudy (anonymous):

Remember this as well `cylindrical seat and four rectangular prism `

OpenStudy (anonymous):

i think i did it wrong i got 244.92

OpenStudy (anonymous):

For which?

OpenStudy (anonymous):

S A

OpenStudy (anonymous):

Well you have the formula can you set it up please, so I know what you're doing?

OpenStudy (anonymous):

SA=(2πrh)+(2πr2)

OpenStudy (anonymous):

Surface area of cylinder: \[2 \pi r h +2 \pi r^2\] Surface area of rectangular prism:\[2lw+2lh+2wh\]

OpenStudy (anonymous):

I mean plug in the numbers and show me

OpenStudy (anonymous):

2 x 3.14 x 3 x 3 + 2 x 3.14 x 3 = 75.36 @iambatman

OpenStudy (anonymous):

thats the cylinder

OpenStudy (anonymous):

the radius half of the diameter, so the radius is 5 not 3.

OpenStudy (anonymous):

`The seat has a diameter of ten inches and a height of 3 inches.`

OpenStudy (anonymous):

oh ok thank u

OpenStudy (anonymous):

Yeah so the surface area of the cylinder is what now?

OpenStudy (anonymous):

You should've got 251.2 for the cylinder surface area\[2 \times 3.14 \times 5 \times 3 + 2 \times 3.14 \times 5^2 = 251.2\]

OpenStudy (anonymous):

Do you understand, did you get the rectangular prism? Make sure to multiply that by 4 as there are 4 of them.

OpenStudy (anonymous):

Afterwards you'll have to add the areas together for the total area